# Interactive Inference: A Neuromorphic Theory of Human-Computer Interaction

### v3

@@ 8. Discussion
8.1. Empirical Study
The above results are a first step towards empirical validation of both our capacity and error models. Results show a good to excellent fit between between predicted and observed models of the task. Driver information processing capacity during deceleration events in a simulated car following task can be measured in real time using the simplified KL divergence of the noise-to-signal ratio of the distance to the car followed, over a time window. Our hypotheses that both distance $S$ and its variance $N$ to the lead car significantly affect the amount of information consumed by the driver during such events were confirmed, with more information consumed at lower $S$s and higher $N$s.
@@ 8.1. Empirical Study
Qualitative ratings by participants were in line with the above. They rated the task as more difficult with higher noise-to-signal ratios (and thus $ID$s), albeit with bipolar ratings of middle $ID$s. We attribute this to participants perceiving distant targets $(S>4.8~{}m)$ as easier to process than close targets with similar signal-to-noise ratios, perhaps due to the 2D nature of our simulator setup. We again note that our results are limited to modeling driver responses to lead car decelerations in a simulated car following task using specific road conditions, car physics, and a sample set of participants.
8.2. Interactive Inference Model and Future Directions
@@
Our results provide good support for the Interactive Inference model, which shows how capacity for performance can be related to error, both in bits. We model average capacity as a logarithmic function of SNR, scaled by a slope that is empirically determined. We can consider this slope to be analogous to the bandwidth in the Shannon-Hartley theorem. Because capacity beyond the limits of this model cannot occur without error, and because both metrics can be converted to bits, one of the predictions from our model is that error can be analysed in view of leftover capacity. This is because the average capacity is, in fact, the maximum capacity. Error subsequently tracks Bayesian surprise as a KL divergence. Figure 6 shows this relationship, with the KL divergence in red fit to some empirically observed error rate (converted to negative log probability) visualized in red dots. Capacity is modeled as a logarithmic function (in blue), which is an empirical fit to Bayesian surprise processed at different SNRs of the task. By subtracting the logarithmic function from this value, we can arrive at a leftover capacity metric (in black). Where this line reaches zero capacity, errors increase and start tracking the KL divergence: the square of the signal-to-noise ratio. The empirical evaluation of this relationship is considered a future direction.
Left-over capacity would correspond directly to the cognitive or sensorimotor load of the task. By correlating empirical results with Heart Rate Variability (HRV) data and other measures of cognitive load evidence for this relationship could be examined. One important topic of investigation is to determine if cognitive load could be measured during tasks, in real time. E.g., users could complete a task modeled via our method while wearing a heart rate monitor. HRV should be lower during moments in the task where there is significant left-over surprise.
The correlation between HRV and cognitive load suggests it may be possible that left-over Bayesian surprise also tracks stress levels in participants. The relationship between our models and experienced stress levels of participants during tasks is therefore an interesting area of future study. E.g., users could complete a task modeled via our method and then fill out a questionnaire about their stress levels. Left-over surprise metrics should correlate with questionnaire results if our hypothesis is correct.
@@
Because our metrics are in bits, rather than time, it should be possible to compare both capacity and error between different tasks. Analyzing different tasks for a signal-to-noise ratio parameter that can be measured, and comparing results from empirical investigations of the relationship between SNR and both capacity and error, would further evidence this assertion. E.g., users could complete different tasks while wearing HRV monitors. Left-over capacity metrics should correlate not just with HRV data, but also between tasks, if our hypothesis is correct. We consider such empirical work a future direction and invite others to help investigate this promising area of research.
9. Conclusions
@@
Neuromorphic Human-Computer Interaction aims to design better user interfaces using strategies that more closely resemble the neurophysiology of the brain. Interactive Inference applies the neuroscience theory of Active Inference to user interface design, postulating that users perform Bayesian inference on current progress and goal distributions to learn tasks and predict their next action. The Bayesian surprise (relative entropy) between goal and progress distributions can be modeled as a square function. However, human capacity to process Bayesian surprise follows a logarithmic function. This explains why errors occur when the user runs out of capacity. We believe our model may allow the quantitative analysis of performance and error in a single framework that allows estimation of the mental load of a task and showed how three basic laws of HCI, Hick’s Law, Fitts’ Law and the Power Law of Practice can be derived from Interactive Inference. We evaluated the use of the model in an empirical study. In our simulated car following task, driver information processing capacity during braking events was linearly related, with good fit, to a logarithmic model of the noise-to-signal ratio of the distance to a lead car, plus 1 (in bits). We found a significant effect of both distance $S$ and variation $N$ on capacity, with a mean driver Index of Performance of almost 4 bits of difficulty (as defined by $log_{2}(N/S+1))$ per bit of information. Participants’ qualitative ratings followed the same trend, with significantly higher ratings of difficulty with higher $N/S$ ratios. We used our model to derive a remaining capacity for the driver, by subtracting the log-linear curve from mean capacity, in bits. We also fit error (collision) rates, with excellent fit, to the square of the signal-to-noise ratio of the distance to the lead car, in bits. We hypothesize that collisions are due to unprocessed information where driver capacity is exceeded. Overall, the empirical results provide support for our Interactive Interference model of HCI. While car-following is just one case study, it provides a framework for conducting further empirical research in Interactive Inference, and evidence that such research is well worth doing. We invite other researchers to test and apply this model to other tasks, to see how well these models predict human performance, and to confirm or refute the general applicability of Neuromorphic Human-Computer Interaction in general, and Interactive Inference in particular.

### v4

@@ Discussion
Empirical Study
The above results are a first step towards empirical validation of both our capacity and error models. Results show a good to excellent fit between between predicted and observed models of the task. Driver information processing capacity during deceleration events in a simulated car following task can be measured in real time using the simplified KL divergence of the noise-to-signal ratio of the distance to the car followed, over a time window. Our hypotheses that both distance $S$ and its variance NN to the lead car significantly affect the amount of information consumed by the driver during such events were confirmed, with more information consumed at lower $S$s and higher NNs.
@@
Qualitative ratings by participants were in line with the above. They rated the task as more difficult with higher noise-to-signal ratios (and thus IDIDs), albeit with bipolar ratings of middle IDIDs. We attribute this to participants perceiving distant targets (S>4.8m)(S>4.8~m) as easier to process than close targets with similar signal-to-noise ratios, perhaps due to the 2D nature of our simulator setup. We again note that our results are limited to modeling driver responses to lead car decelerations in a simulated car following task using specific road conditions, car physics, and a sample set of participants.
Interactive Inference Model and Future Directions
@@
Our results provide good support for the Interactive Inference model, which shows how capacity for performance can be related to error, both in bits. We model average capacity as a logarithmic function of SNR, scaled by a slope that is empirically determined. We can consider this slope to be analogous to the bandwidth in the Shannon-Hartley theorem. Because capacity beyond the limits of this model cannot occur without error, and because both metrics can be converted to bits, one of the predictions from our model is that error can be analysed in view of leftover capacity. This is because the average capacity is, in fact, the maximum capacity without error. Error subsequently tracks Bayesian surprise as a KL divergence. Figure 6 shows this relationship, with the KL divergence in red fit to some empirically observed error rate (converted to negative log probability) visualized in red dots. Capacity is modeled as a logarithmic function (in blue), which is an empirical fit to Bayesian surprise processed at different SNRs of the task. By subtracting the logarithmic function from this value, we can arrive at a leftover capacity metric (in black). Where this line reaches zero capacity, errors increase and start tracking the KL divergence: the square of the signal-to-noise ratio. The empirical evaluation of this relationship is considered a future direction.
@@
Left-over capacity would correspond directly to the cognitive or sensorimotor load of the task. By correlating empirical results with Heart Rate Variability (HRV) data and other measures of cognitive load evidence for this relationship could be examined. One important topic of investigation is to determine if cognitive load could be measured during tasks, in real time. E.g., users could complete a task modeled via our method while wearing a heart rate monitor. HRV should be lower during moments in the task where there is significant left-over surprise. The correlation between HRV and cognitive load suggests it may be possible that left-over Bayesian surprise also tracks stress levels in participants. The relationship between our models and experienced stress levels of participants during tasks is therefore an interesting area of future study. E.g., users could complete a task modeled via our method and then fill out a questionnaire about their stress levels. Left-over surprise metrics should correlate with questionnaire results if our hypothesis is correct. Because our metrics are in bits, rather than time, it should be possible to compare both capacity and error between different tasks. Analyzing different tasks for a signal-to-noise ratio parameter that can be measured, and comparing results from empirical investigations of the relationship between SNR and both capacity and error, would further evidence this assertion. E.g., users could complete different tasks while wearing HRV monitors. Left-over capacity metrics should correlate not just with HRV data, but also between tasks, if our hypothesis is correct. We consider such empirical work a future direction and invite others to help investigate this promising area of research.
Limitations
This work presents an exploratory framework inspired by Active Inference principles that does not require implementation of its full mathematical machinery (such as model evidence and expected free energy). Our use of the term Bayesian surprise refers specifically to the KL divergence between our operationally-defined goal and progress distributions, not a formal construct from Active Inference theory. In addition, the connections we draw to Fitts’ Law, Hick’s Law, and the Power Law are observational. Finally, our empirical validation is limited to a single driving task, and further research is needed to test the generalization of this approach to other interaction contexts.
Conclusions
@@
Neuromorphic Human-Computer Interaction aims to improve user interfaces by exploring how concepts from neuroscience might inform design. Interactive Inference offers a simplified interpretation of the neuroscience theory of Active Inference in user interface design. It postulates that users perform Bayesian inference on current progress and goal distributions to learn tasks and predict their next action. The Bayesian surprise (relative entropy) between goal and progress distributions can be modeled as a square function. However, human capacity to process Bayesian surprise follows a logarithmic function. This explains why errors occur when the user runs out of capacity. We believe our model may allow the quantitative analysis of performance and error in a single framework that allows estimation of the mental load of a task and showed how three basic laws of HCI, Hick’s Law, Fitts’ Law and the Power Law of Practice can be derived from Interactive Inference. We evaluated the use of the model in an empirical study. In our simulated car following task, driver information processing capacity during braking events was linearly related, with good fit, to a logarithmic model of the noise-to-signal ratio of the distance to a lead car, plus 1 (in bits). We found a significant effect of both distance $S$ and variation NN on capacity, with a mean driver Index of Performance of almost 4 bits of difficulty (as defined by $log_{2}(N/S+1))$ per bit of information. Participants’ qualitative ratings followed the same trend, with significantly higher ratings of difficulty with higher $N/S$ ratios. We used our model to derive a remaining capacity for the driver, by subtracting the log-linear curve from mean capacity, in bits. We also fit error (collision) rates, with excellent fit, to the square of the signal-to-noise ratio of the distance to the lead car, in bits. We hypothesize that collisions are due to unprocessed information where driver capacity is exceeded. Overall, the empirical results provide support for our Interactive Interference model of HCI. While car-following is just one case study, it provides a framework for conducting further empirical research in Interactive Inference, and evidence that such research is well worth doing. We invite other researchers to test and apply this model to other tasks, to see how well these models predict human performance, and to confirm or refute the general applicability of Neuromorphic Human-Computer Interaction in general, and Interactive Inference in particular.